A rope 33 ft long is stretched between a tree and a car. A man pulls with a force of 3 Lbs at right angles to and at the middle point of the rope, and move this point 20 inches. Assuming no stretching of the rope, what is the tension in the rope at the final position?


1
Expert's answer
2021-05-26T10:31:47-0400

The tension is

"T=\frac{F}{2\sin\theta}=\frac{F}{2(2h/L)}=\frac{FL}{4h},\\\space\\
T=\frac{3\cdot33}{4\cdot 1.67}=14.8\text{ lbs}."